The concentrated sulfuric acid that we use in the laboratory is 98.0 % H2SO4 by mass. Calculate the molality and molarity of the solution. The density of the solution is 1.83 g/mL.

1
Expert's answer
2022-03-02T03:09:51-0500

Mass of H2 SO4 = 98%



Means 98gm of H2 SO4 solute are present per 100gm of solution.



Hence mass of H2SO4 solute = 98gm.



Molar mass of H2SO4 solute =molar mass = (1.008 g/mol × 2) + (32.07 g/mol × 1) + (16.00 g/mol × 4)



Molar mass = 98.086 g/mol



Mass of solution = volume x density


= 1000 x 1.83 = 1830 g



Mass of H2SO4 = 98.0/100 x 1830 = 1793.4 g



Moles of H2SO4 = mass/molar mass of H2SO4


= 1793.4/98.08


= 18.3 mol



Molarity = No. Of moles of H2SO4 solute / Volume of solution in l


=18.3/1


=18.3M



Therefore, the molarity of


solution H2SO4 is 18.3 M



Molality = moles of H2SO4/mass of water in kg



= 18.3/0.0366



= 500 mol/kg = 500m